OpenStudy (anonymous):

Consider a random variable X such that represents the throw of a fair die. Let Y be a random variable where P(Y=1) = (Your birth month)/15; P(Y=2), P(Y=3), P(Y=4), P(Y=5) and P(Y=6) are equally probable. The sample space for each variable is {1,..., 6} Fill in columns 2 and 3 in the following table where each entry is the probability of the random variable having the indicated outcome. Table: n | P(X=n) | P(Y=n) 1 2 3 4 5 6 (Note: Birth month = 3)

OpenStudy (anonymous):

go on mathway.com maybe it could help you w/ your question? :) Let me know if it helped

OpenStudy (anonymous):

It didn't help :( But it's a very interesting site!

OpenStudy (anonymous):

What's the probability that you roll a 1 with a die if there are 6 possible faces?

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

Right. This is the same as saying \(P(X=1)=\dfrac{1}{6}\). Any roll has an equal outcome, so \(P(X=n)=\dfrac{1}{6}\) for all \(n\).

OpenStudy (anonymous):

Yup :) I understand that I'm mostly confused about the whole P(Y=1) = (Your birth month)/15. Does this mean P(Y=1) =3/15 then the whole thing over 5?

OpenStudy (phi):

***Does this mean P(Y=1) =3/15*** this part is true. or more simply, P(Y=1)= 1/ 5 to find P(Y= n) where n > 1, we use this info: 1. sum of all the P's must be 1 P(Y=1) + P(Y=2) + P(Y=3) + P(Y=4) + P(Y=5) + P(Y=6) =1 2. P(Y=2), P(Y=3), P(Y=4), P(Y=5) and P(Y=6) are equally probable. if we call P(Y=2) =x (i.e. the unknown probability) then we can say, using 1. , P(Y=1) + P(Y=2) + P(Y=3) + P(Y=4) + P(Y=5) + P(Y=6) =1 1 / 5 + 5x = 1 or 0.2 + 5x = 1 add -0.2 to both sides 5x = 1-0.2 or 5x = 0.8 x = 0.8/5 = 0.16 now you can fill in the P(Y) column